# Why Transformers Need Adam: A Hessian Perspective

### v1

@@ Abstract
SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear. In this work, we provide an explanation of SGD’s failure on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity. To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists. Our initial theoretical analysis indicates that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. The failure could be rescued if we could assign different learning rates across blocks, as designed in Adam.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have become a major workhorse behind much remarkable progress in AI development (e.g., (Achiam et al., 2023)). However, the understanding of Transformer training remains limited. For instance, Transformer training largely relies on Adam(W) (Kingma and Ba, 2014; Loshchilov and Hutter, 2017). In contrast, Stochastic Gradient Descent with momentum (SGD)  [^1] We introduce the update rules of Adam(W) and SGD in Appendix B.1., which is the de-facto optimizer for convolution neural networks (LeCun et al., 1998) (CNNs), performs poorly on Transformers (see Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in Appendix D as evidence). Yet, the cause of failure remains unclear. It is an intriguing question why SGD fails on Transformers. First, from a theoretical perspective, this topic helps reveal more properties of the Transformer training loss landscape, which helps boost understanding of training objectives and current optimizers. Second, from a computational perspective, Adam requires heavier memory storage than SGD. Compared with SGD, Adam needs to store an additional second-order momentum, which takes at least the same memory of the model size. The memory consumption of Adam has become a major overhead in large-scale transformer training (Rasley et al., 2020). Identifying the cause of failure may help researchers design more efficient optimizers.
@@
In this work, we explore the failure mode of SGD through the lens of Hessian. We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum as an example). By theory, the full Hessian spectrum largely determines the behavior of gradient methods (Nesterov, 2013; Goh, 2017; Sun, 2019; Goujaud et al., 2022), so we suspect it can also help uncover the reasons behind SGD failure on Transformers. Using numerical linear algebra tools (Bai et al., 1996), we empirically compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). Unfortunately, in Figure LABEL:fig_full_spectrum, we find that the spectra on CNNs and Transformers are often largely similar despite the different optimizer behaviors on these tasks. As such, we have not identified critical phenomena in the full Hessian spectra associated with SGD’s failure on Transformers. To reveal the cause of failure, we need a more fine-grained investigation into Hessian.
@@
What else could make SGD fail on Transformers, but not on CNNs? By dissecting the structures of CNNs and Transformers, we observe that CNN architectures are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the interleaved stacking of disparate parameter blocks (e.g. Query, Key, Value blocks in attention and MLP layers). We suspect that these design differences may lead to different properties in optimization. Intuitively, disparate parameter blocks would contribute differently to the total loss. So each block might prefer a distinct treatment by optimizers, which can be provided by Adam but not by SGD. This motivates us to investigate the Hessian spectrum of each parameter block, abbreviated as the blockwise Hessian spectrum.
By inspecting the blockwise Hessian spectrum, we find a possible explanation why SGD fails on Transformers. We find that Transformers are “heterogeneous": for all Transformers we checked, Hessian spectra of different parameter blocks are dramatically different. We then verify that such heterogeneity indeed hampers SGD via extensive experiments on various Transformers, CNNs, MLPs , and quadratic problems. Our contributions are summarized as follows.
@@
  • Exploring the failure mode of SGD on transformers. We provide an explanation why SGD largely lags behind Adam on Transformers through the lens of blockwise Hessian spectrum. This explanation consists of two parts. First, we identify a phenomenon called “block heterogeneity", which states that Hessian spectra are dramatically different across parameter blocks. We find this phenomenon prevalent on all Transformers in our experiments, but not on CNNs. Second, we verify that block heterogeneity hampers SGD. More specifically, on various kinds of Transformers, CNNs, MLPs, we find that SGD always performs badly on problems with block heterogeneity but works well otherwise.
@@
  • Theoretical results on quadratic models. We build various convex quadratic problems, some with block heterogeneity and others without block heterogeneity. We find that GD is significantly worse than Adam on problems with block heterogeneity, but performs similarly to Adam otherwise. We provide a theoretical result which shows that GD indeed works badly on quadratic problems with block heterogeneity. Our analysis suggests that one important reason for (S)GD failure is that it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. The failure could be saved if we assign different learning rates across blocks like Adam.
@@ 2.1 Problem Settings
We denote the full-batch training loss as $\mathcal{L}(w)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(w)$, where $n$ is the number of minibatches, $\mathcal{L}_{i}(w)$ is the loss of $i$-th minibatch and $w\in\mathbb{R}^{d}$ is the neural network parameters. We denote the gradient and Hessian of the training loss w.r.t. neural network parameters as $\nabla\mathcal{L}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}(w)\in\mathbb{R}^{d\times d}$, respectively. We use $\nabla\mathcal{L}_{i}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}_{i}(w)\in\mathbb{R}^{d\times d}$ to denote the $i$-th minibatch counterparts. We use $[d]$ to denote the index set $\{1,2,\cdots,d\}$. Given an arbitrary partition $\{\mathcal{D}_{l}\}^{L}_{l=1}$ over $[d]$ with $d_{l}\triangleq|\mathcal{D}_{l}|$, we can split $w$ into $L$ parameter blocks $\{w_{l}\}_{l=1}^{L}$, where $w_{l}=\mathbb{R}^{d_{l}}$ consists of parameters with indexes in the $l$-th block $\mathcal{D}_{l}$. We denote $[\nabla^{2}\mathcal{L}(w)]_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ as the Hessian of $l$-th parameter-block $w_{l}$, where $[\nabla^{2}\mathcal{L}(w)]_{l,i,j}=\frac{\partial^{2}}{\partial_{w_{l,i}}\partial_{w_{l,j}}}\mathcal{L}(w_{l})$. Note that $[\nabla^{2}\mathcal{L}(w)]_{l}$ is the $l$-th principle block sub-matrix of $\nabla^{2}\mathcal{L}(w)$.
Preliminaries.
Hessian of large-scale NNs are intractable to obtain due to the numerical difficulties of computing and storing them. Fortunately, there are extensive numerical linear algebra methods to approximate the Hessian spectrum without calculating the Hessian explicitly. Among which, the Stochastic Lanczos Quadrature method (SLQ) (Bai et al., 1996) is known to be one of the oldest yet the most powerful methods. SLQ is a combination of the stochastic trace estimator (Hutchinson, 1989), the Gaussian quadrature (Golub and Meurant, 2009; Epperson, 2013), and the Lanczos algorithm (Lanczos, 1950). In short, SLQ takes a handful of Hessian-vector products as input, where the vectors are sampled from the normalized Rademacher distribution. It returns a smooth curve on $\mathbb{R}$ that approximates the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example). Since the derivation of SLQ is quite involved, we defer the detailed description of SLQ to Appendix B.2. The algorithmic form of SLQ can be seen in Algorithm 5 in the Appendix. We provide running time analysis and detailed setup of SLQ in Appendix C.1.
In this work, we apply SLQ to various CNNs and Transformers on different datasets. As for CNNs, we study CNNs including ResNet18 (11M) and VGG16 (138M) on ImageNet (He et al., 2016; Simonyan and Zisserman, 2014). As for Transformers, we study ViT-base (86M) on ImageNet (Dosovitskiy et al., 2020), BERT (40M) on Cornell Movie-Dialogs Corpus (Devlin et al., 2018), and GPT2 (125M) on Openwebtext (Radford et al., 2019). For a fair comparison with ResNet18, we also consider smaller models like GPT2-nano [^2] https://github.com/karpathy/nanoGPT/ (11M) on English corpus. The largest model we conduct is GPT2 with $125$ million parameters. Detailed experimental setups are shown in Appendix C.1. For all these Transformer-based tasks, SGD performs significantly worse than Adam; while for CNN-based takes, Adam and SGD are similar. The performance comparison is shown in Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in the Appendix. We estimate the spectrum of (1) the full Hessian $\nabla^{2}\mathcal{L}(w)$ and (2) the Hessian of all parameter blocks $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$. For (2), we split $w$ according to the default partition in PyTorch implementation, e.g., Embedding layer, Query in 1st attention layer, Key in 1st attention layer, Value in 1st attention layer, etc..
@@
Note that the term “block" is a bit different from the term “layer". For instance, Query and Key could be in the same layer, but they are different parameter blocks. In the following contexts, we refer to “full Hessian spectrum" as the estimated spectrum of $\nabla^{2}\mathcal{L}(w)$ and “blockwise Hessian spectrum" as the estimated spectrum of $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
Motivated by classical theory, we start by investigating the full Hessian spectrum of Transformers. We choose to study the full Hessian spectrum for two reasons. First, as mentioned in Section 1, the full Hessian spectrum is known to largely determine the behavior of gradient methods. For instance, the range of the spectrum will affect the convergence rate (Nesterov, 2013). Second, the full Hessian spectrum has been shown to be informative in explaining multiple phenomena in neural nets (NNs), e.g., how BatchNorm accelerates the training (Ghorbani et al., 2019). Due to these reasons, we hypothesize that the full Hessian spectrum can also help uncover the reasons behind SGD failure on Transformers.
We compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). The results are shown in Figure LABEL:fig_full_spectrum. Unfortunately, our experiments show that the full Hessian spectrum might not be enough to explain the SGD failure on Transformers. We explain as follows. The primary information of the spectrum lies in its (1) dispersion, (2) shape, and (3) evolution along training. As for (1), we find that eigenvalues are dispersed very similarly and we do not observe any notably large eigenvalues in the spectra of Transformers. So (1) does not seem to be related to the failure of SGD. We further inspect (2) and (3). For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we find similar phenomena: the shape of the spectrum is approximately symmetrical around 0 at initialization. As training proceeds, the majority of large negative eigenvalues disappear and the shape evolves into a combination of a “bulk" and some “outliers". Since the spectral shape and evolution are quite similar on Transformers and CNNs, they cannot explain the failure of SGD on Transformers, either. In summary, we have not identified critical phenomena in the full Hessian spectra linked to SGD’s failure on Transformers.
@@ 2.3 Main Findings Through Blockwise Hessian Spectra
We have shown that the full Hessian spectrum seems not sufficient for uncovering the reason why SGD fails on Transformers. What else could make SGD fail on Transformers, but not on CNNs? We find two important features that are overlooked in the full Hessian spectrum analysis above.
@@
  • (1) The Hessian structure is overlooked. There are numerical results showing that the Hessians are close to block-diagonal matrices on MLPs (Collobert, 2004; Roux et al., 2007; Martens and Grosse, 2015). We restate their findings in Appendix C.2. Further, (Collobert, 2004, Section 7) theoretically shows that the block diagonal structure comes from (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss. Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principle block correspond to the parameters of each block in the Transformer. These results suggest that near block-diagonal Hessian might be common in NNs.
@@
  • (2) The build-up rules of Transformers is overlooked. Transformers are built upon stacking of disparate parameter blocks, e.g. Query, Key, Value blocks in attention, and MLP layers. This is largely different from CNNs, which are constructed by the repetitive stacking of similar parameter blocks (convolution layers). Since each parameter blocks in Transformers are designed differently, they might have different properties for optimization, which might further affect the optimizer behavior.
@@
Combining (1) and (2), we hypothesize that the blockwise Hessian spectrum might provide some insights on distinguishing CNNs and Transformers. What extra information can be contained in the blockwise spectrum but not in the full spectrum? By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian. For block diagonal matrices, blockwise Hessian encodes the location of eigenvalues, i.e., which block does an eigenvalue (of the full Hessian) lie in, while the full spectrum ignores this. In the following, we study blockwise Hessian spectra on various models. We will show that blockwise spectra indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
@@
We here demonstrate the shape of blockwise spectra in CNNs and Transformers. We sample four blocks per model and present the spectra in Figure LABEL:fig_blockwise_spectrum. In Transformers, the Hessian spectra of embedding, attention, and MLP blocks are largely different. In contrast, in CNNs, the spectra of convolution layers are similar. We further verify this observation for the rest of the parameter blocks. We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs. We show the results in Figure LABEL:fig_blockwise_heatmap.
@@
The results in Figure LABEL:fig_blockwise_heatmap showed an interesting phenomenon: for all the Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". As a control group experiment, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneous Hessian in a CNN comes from the repetitively similar convolution layers, while the block-heterogeneous Hessian in a Transformer stems from the interleaved stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with the failure of SGD on Transformers.
2.4 SGD Fails on Simple MLPs with Block Heterogeneity
We now further link block heterogeneity to the failure of SGD. We construct a simple example without attention layers but has block heterogeneity and find that SGD still fails. We consider a 4-layer MLP on MNIST. Using standard MLP, we find that the Hessian exhibits a homogeneous block structure and SGD converges as fast as Adam. We then scale the output of each layer by a constant $c$. We define $c$ as the degree of heterogeneity since larger $c$ would bring more heterogeneity to the Hessian (see Table 3 in the Appendix). As shown in Figure LABEL:fig_mlp_gap, we find that the gap between Adam and SGD increases with the degree of heterogeneity. This example shows that attention might not be the root cause of SGD failure, but the block heterogeneity (brought by attention) might be.

### v2

@@ Abstract
SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear. In this work, we provide an explanation of SGD’s bad performance on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity. To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists. Our initial theoretical analysis indicates that SGD performs poorly because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks. This limitation could be ameliorated if we use coordinate-wise learning rates, as designed in Adam. [^1] Our code is available at https://github.com/zyushun/hessian-spectrum.
{NoHyper}*: Correspondence author.
@@ 1 Introduction
Transformers [84] have become a major workhorse behind much remarkable progress in AI development (e.g., [1]). However, the understanding of Transformer training remains limited. For instance, Transformer training largely relies on the Adam optimizer [45, 56]. In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs poorly on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum). Yet, the reasons behind this poor performance remain unclear. Understanding why SGD performs badly on Transformers is an intriguing question. First, from a theoretical perspective, this can help us better understand the training of Transformers and more generally, neural networks. Second, from a computational perspective, the understanding may inspire the design of better algorithms for training neural networks.
@@
In this work, we explore the cause for SGD’s deficiency through the lens of Hessian. We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum). By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s bad performance on Transformers. Using tools from numerical linear algebra [7], we empirically compare the full spectra of CNNs (where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). Unfortunately, as shown in Figure LABEL:fig_full_spectrum, the spectra for CNNs and Transformers are often largely similar despite the different optimizer behaviors. As such, we have not identified critical features in the full Hessian spectra associated with SGD’s bad performance on Transformers. To reveal the cause, a more fine-grained investigation into the Hessian is needed.
@@
What would cause SGD to perform badly on Transformers but not on CNNs? By dissecting the structures of CNNs and Transformers, we notice that CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the non-sequential stacking of disparate parameter blocks (e.g. Query, Key, Value, Output projection blocks in attention and MLP layers). We hypothesize that these architectural differences might lead to different optimization properties. Intuitively, disparate parameter blocks contribute differently to the overall loss. So each block might benefit from a specialized treatment by optimizers, a flexibility offered by Adam but not by SGD. This observation motivates us to investigate the Hessian spectrum of each parameter block, which we refer to as the blockwise Hessian spectrum.
By inspecting the blockwise Hessian spectrum, we discover a possible explanation for SGD’s bad performance: the “heterogeneity” inherent in Transformers. We provide both empirical and theoretical evidence to support this explanation. Our contributions can be summarized as follows:
@@
  • Failure mode of SGD on Transformers. We explain why SGD underperforms Adam on Transformers by examining the blockwise Hessian spectrum. First, we identify a phenomenon called ’block heterogeneity’, which refers to the large differences in the Hessian spectra across parameter blocks. This block heterogeneity is observed in all examined Transformers but not in CNNs. Second, we verify that block heterogeneity hinders SGD. Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs poorly on problems with block heterogeneity but works well otherwise.
@@
  • Theoretical results on quadratic models. We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) performs significantly worse than Adam on problems with block heterogeneity but comparably otherwise. Our theoretical analysis shows that GD struggles on quadratic problems with block heterogeneity, suggesting that using a single learning rate for all blocks is a key reason for GD’s bad performance. The deficiency can be mitigated by assigning different learning rates across blocks, as Adam does.
We emphasize that we do not claim block heterogeneity is the only cause for SGD’s bad performance, but just that it is at least one important cause. We verify, both empirically and theoretically, that SGD performs poorly when block heterogeneity is present.
@@ 2.1 Problem Settings
Notations. We denote the training loss as $\mathcal{L}(w)$, where $w\in\mathbb{R}^{d}$ is the neural network parameters. We denote the gradient and Hessian of the training loss w.r.t. neural network parameters as $\nabla\mathcal{L}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}(w)\in\mathbb{R}^{d\times d}$, respectively. We use $[d]$ to denote the index set $\{1,2,\cdots,d\}$. Given an arbitrary partition $\{\mathcal{D}_{l}\}^{L}_{l=1}$ over $[d]$ with $d_{l}\triangleq|\mathcal{D}_{l}|$, we can split $w$ into $L$ parameter blocks $\{w_{l}\}_{l=1}^{L}$, where $w_{l}=\mathbb{R}^{d_{l}}$ consists of parameters with indexes in the $l$-th block $\mathcal{D}_{l}$. We denote $[\nabla^{2}\mathcal{L}(w)]_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ as the Hessian of $l$-th parameter-block $w_{l}$, where $[\nabla^{2}\mathcal{L}(w)]_{l,i,j}=\frac{\partial^{2}}{\partial_{w_{l,i}}\partial_{w_{l,j}}}\mathcal{L}(w_{l})$. Note that $[\nabla^{2}\mathcal{L}(w)]_{l}$ is the $l$-th principal block sub-matrix of $\nabla^{2}\mathcal{L}(w)$.
Setups. Hessian of large-scale NNs are intractable to compute and store. In this work, we apply a numerical tool called Stochastic Lanczos Quadrature method (SLQ) [7] to approximate the Hessian spectrum. SQL uses a smooth curve on $\mathbb{R}$ to approximate the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example). A detailed introduction to SLQ is provided in Appendix C.2. All experimental setups in this section is shown in Appendix D. We focus primarily on the following models/tasks.
  • CNNs. We study ResNet18 (11M) and VGG16 (138M) on ImageNet [40, 79]. On these tasks, SGD performs on par with Adam. See Figure LABEL:fig:cv_figure in Appendix B for the evidence.
  • Transformers. We study Transformer with various scales and modalities, including GPT2 (125M) on OpenWebText [71]; ViT-base (86M) on ImageNet [27]; BERT (40M) on Cornell Movie-Dialogs Corpus [25]; GPT2-nano [^3] https://github.com/karpathy/nanoGPT/ (11M) on English corpus. On these tasks, SGD performs significantly worse than Adam. See Figure LABEL:fig:nlp_figure in Appendix B for the evidence.
@@
For each model, we estimated (1) the full Hessian spectrum $\nabla^{2}\mathcal{L}(w)$, and (2) the blockwise Hessian spectrum $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$. For the latter, we split $w$ according to the default partition in PyTorch implementation, e.g., Embedding layer, Query in each attention layer, Key in each attention layer, Value in each attention layer, etc. Note that the term “block" differs from the term “layer". For instance, Query and Key can reside in the same layer but are different parameter blocks.
@@ 2.2 Full Hessian Spectrum Is Not Informative Enough
We study the full Hessian spectrum of Transformers for two reasons. First, as stated in Section 1, the Hessian spectrum significantly influences the behavior of gradient methods [63]. Second, previous research shows that the Hessian spectrum provides insights into neural network phenomena, like BatchNorm’s effect on training speed [32]. Therefore, we hypothesize that the Hessian spectrum may also explain SGD’s bad performance on Transformers.
We compare the full Hessian spectra of CNNs (where SGD works well) and those of Transformers (where SGD struggles), as shown in Figure LABEL:fig_full_spectrum. Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain why SGD struggles on Transformers. We elaborate as follows. The primary information in the spectrum lies in its (A) dispersion, (B) shape, and (C) evolution during training. Regarding (A), we observe that the eigenvalues are dispersed similarly across different models, with no notably large outlier for Transformers. Thus, dispersion does not seem to be related to the bad performance of SGD. We further investigate (B) and (C). For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we observe similar phenomena: the spectrum’s shape is approximately symmetrical around 0 at initialization. As training proceeds, the majority of negative eigenvalues disappear, and the shape evolves into a combination of a “bulk” and some “outliers”. Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain the incompetence of SGD on Transformers either. In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to SGD’s bad performance on Transformers.
@@ 2.3 Main Findings Through Blockwise Hessian Spectra
What other factors could cause SGD to struggle on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
  • 1.
@@
    Hessian structure. Existing literatures showed that the Hessians of MLPs are close to block-diagonal matrices [18, 73, 60]. We restate their findings in Appendix B. Further, [18, Section 7] theoretically attributes the block diagonal structure to (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss. Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principal block correspond to the parameters of each block in the Transformer. These results suggest that near block-diagonal Hessian might be common in NNs.
  • 2.
@@
    Build-up rules of Transformers. As shown in Figure LABEL:fig_blockwise_spectrum, CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers). In contrast, Transformers consist of disparate parameter blocks, e.g. Query, Key, Value in attention, and MLP layers. Further, these blocks are stacked in a non-sequential manner. Due to the different designs of these blocks, they may have different properties for optimization, which can further affect the optimizer behavior.
@@
Combining these together, we hypothesize that the blockwise Hessian spectrum, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$, might provide additional insights. What extra information can be contained in the blockwise spectrum but not in the full spectrum? By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian. We note that Transformers are observed to have a near block-diagonal Hessian. For block diagonal matrices, blockwise spectra encode the location of eigenvalues, i.e., which block an eigenvalue (of the full matrix) resides in. In contrast, the full spectrum overlooks this location information. In the following, we study the blockwise Hessian spectra of various models and show that they indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
@@
We here demonstrate the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer). We sample four blocks for each model and present the spectra in Figure LABEL:fig_blockwise_spectrum. In BERT, the Hessian spectra of embedding, attention, and MLP blocks are largely different. In contrast, in ResNet, the spectra of convolution layers are similar. We further verify this observation for the rest of the parameter blocks. We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs and show the results in Figure LABEL:fig_blockwise_heatmap.
@@
The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with SGD’s bad performance on Transformers.
2.4 SGD Performs Badly on Various Tasks with Block Heterogeneity
Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD performs badly on Transformers. (2) Transformers have block-heterogeneity. Now we further link block heterogeneity to SGD’s bad performance on non-Transformer models. This would directly establish a connection between block heterogeneity and SGD’s bad performance, without going through Transformers or attention blocks as an intermediary. We consider one man-made example and one real-world example.
Example 1: A man-made MLP. We consider a 4-layer MLP on MNIST and change the degree of heterogeneity by scaling each layer by constant $c$. Figure LABEL:fig_mlp_gap (a) shows SGD fails as heterogeneity grows.
Example 2: MLP-mixer. We consider MLP-mixer [82], a famous all-MLP architecture that outperforms CNNs and ViTs on some vision tasks. Figure LABEL:fig_mlp_gap (b) (c) shows that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
2.5 Reduced Block Heterogeneity in Pre-trained Transformers
We remark that different Transformers exhibit different levels of block heterogeneity. Although all examined Transformers show strong block heterogeneity, we find that this heterogeneity can be mitigated, resulting in less performance deterioration for SGD. As illustrated in Figure LABEL:fig_sft_gap, pre-trained GPT2 on SFT tasks can exhibit less block heterogeneity compared to pre-training GPT2 from scratch (Figure LABEL:fig_blockwise_heatmap (f)). In this case, although SGD is still slower than Adam, it achieves a similar loss at convergence. Compared with training GPT2 from scratch (Figure LABEL:fig:nlp_figure (d) in Appendix B), the performance gap between SGD and Adam is significantly narrowed down. These findings suggest that the heterogeneity induced by architectural design can be alleviated by selecting “good” weights. This partly explains why simpler methods like SGD and even its zeroth-order version can still be effective for fine-tuning language models, albeit with slower convergence [58, 59].